OpenStudy (anonymous):

Recall the equation for projectiles is y = –16x2 + vx + s, where y is the object’s current height (in feet) at any time after it is projected, x is the time (in seconds) the object has been in the air, v is the object’s starting upward velocity, and s is the object’s starting height.

OpenStudy (blurbendy):

what is the question

OpenStudy (anonymous):

use graphing technology to complete this activity since many of the solutions will most likely involve irrational numbers. @blurbendy

OpenStudy (blurbendy):

Okay, is there a question where we can use the formula y = -16x^2 + vx + s, or are we just supposed to graph the formula on its own?

OpenStudy (anonymous):

were just supposed to graph it on our own

OpenStudy (blurbendy):

Okay, do you have a basic graph Cartesian graph drawn in front of you (a graph with an x-axis and a y-axis)?

OpenStudy (anonymous):

yes

OpenStudy (blurbendy):

Okay, we need starting values for x, v, and s. Let's start with x = 1, v = 2, and s = 3. What does the equation become when you substitute these values into those variables?

OpenStudy (anonymous):

y=-16x^2+2(1)+3

OpenStudy (blurbendy):

close y = -16(1)^2 = -16 So, y = -16+2 +5 y = -9

OpenStudy (blurbendy):

So, on your graph, you can put a point at (1, -9)

OpenStudy (blurbendy):

Don't worry about s or v since they are constants

OpenStudy (blurbendy):

Keep putting in values for x and see what you get for y (you should see a pattern)

OpenStudy (anonymous):

i think i understand now, omg thank you so much!!

OpenStudy (blurbendy):

Any time!